OpenStudy (freckles):

\[\text{ Suppose } f \text{ is an analytic function of the complex variable } z=x+i y \text{ given by } \\ f(z)=(2x+3y)+i g(x,y) \\ \text{ where} g(x,y) \text{ is a real-valued function } \\ \text{ of the real variables } x \text{ and } y . \\ \text{ If } g(2,3)=1\text{, then }g(7,3)=? . \]

OpenStudy (freckles):

\[f(2+3i)=4+9+ig(2,3) \\ f(2+3i)=13+i g(2,3) \\ f(2+3i)=13+i \\ f(7+3i)=14+9+ig(7,3) \\ f(7+3i)=23+i g(7,3) \\ \] this as far as I can get so far

OpenStudy (freckles):

There are choices A) -14 B) -9 C) 0 D) 11 E) 18 This question comes from the mathematics gre practice exam.

OpenStudy (xapproachesinfinity):

hmm interesting, i have some admiration for complex though i have no skills in such things yet lol

OpenStudy (anonymous):

I believe it might have something to do with cancelling the complex part.

OpenStudy (xapproachesinfinity):

i have a feeling that it is not that hard just kind of a trick

OpenStudy (freckles):

was kind of thinking that too

OpenStudy (freckles):

somehow we have to show it is -14

OpenStudy (freckles):

maybe and I could be totally crazy we could do something with magnitudes of complex numbers

OpenStudy (freckles):

\[|g(2,3)|=1 \\ |f(2,3)|=\sqrt{170} \\ |g(7,3)|=\sqrt{a^2} \text{ if } g(7,3)=a \\ |f(7,3)|=\sqrt{23^2+g^2(7,3)}\]

OpenStudy (freckles):

\[f^2(7,3)=23^2+a^2 \]

OpenStudy (xapproachesinfinity):

hmm

OpenStudy (freckles):

still playing

OpenStudy (xapproachesinfinity):

you -14 was the answer! did you just tried it directly?

OpenStudy (xapproachesinfinity):

said*

OpenStudy (freckles):

no I looked at the answers

OpenStudy (xapproachesinfinity):

oh ok

OpenStudy (freckles):

how did you know -14 was right?

OpenStudy (xapproachesinfinity):

no i just saw your reply above mentioning that

OpenStudy (freckles):

I wonder what analytic means

OpenStudy (freckles):

looking up...

OpenStudy (xapproachesinfinity):

isn't that related to graphs ?

OpenStudy (xapproachesinfinity):

don't know why wolfram ignores the imaginary stuff for such functions

OpenStudy (freckles):

\[f(x,y)=u(x,y)+i v(x,y) \\ \frac{\partial u}{ \partial x}=\frac{\partial v}{ \partial y} \\ \frac{\partial u}{ \partial y}=-\frac{ \partial v }{\partial x} \\ 2=g_y \\ 3=-g_x\]

OpenStudy (anonymous):

Analytic functions refers to something else. One important property of analytic functions are that they are indefinitely differentiable.

OpenStudy (xapproachesinfinity):

you know it crossed my mind that we need to use calculus lol

OpenStudy (freckles):

\[g(x,y)=2y+C(x) \\ -g(x,y)=3x+K(y) \\ \\ 0=2y+3x+C(x)+K(y) \] where C(x) is a function of x and K(y) is a function of y

OpenStudy (xapproachesinfinity):

so any analytic function has no differentiability problem is that what you are saying?

OpenStudy (freckles):

\[2g(x,y)=2y-3x+C(x)-K(y) \\ 2g(2,3)=2(3)-3(2)+C(2)-K(3) \\ g(2,3)=C(2)-K(3)\]

OpenStudy (freckles):

err...

OpenStudy (xapproachesinfinity):

oh i see! after checking the theorem in that pdf

OpenStudy (freckles):

lol still trying to play with this fungus

OpenStudy (freckles):

and that is equal to 1

OpenStudy (freckles):

\[1=C(2)-K(3)\]

OpenStudy (freckles):

\[2g(7,3)=2(3)+3(7)+C(7)-K(3)\]

OpenStudy (freckles):

\[2g(7,3)=27+C(7)-K(3)\]

OpenStudy (freckles):

\[2g(7,3)-1=27+C(7)-C(2)\]

OpenStudy (anonymous):

Oh yea, use the Cauchy–Riemann equations to solve this.

OpenStudy (freckles):

\[2g(7,3)=28+C(7)-C(2) \\ g(7,3)=14+\frac{C(7)-C(2)}{2}\]

OpenStudy (freckles):

oops i made a type-o somewhere above

OpenStudy (freckles):

\[g(x,y)=2y+C(x) \\ -g(x,y)=3x+K(y) \\ \\ 0=2y+3x+C(x)+K(y)\] \[2g(x,y)=2y-3x+C(x)-K(y) \\ 2g(2,3)=2(3)-3(2)+C(2)-K(3) \\ g(2,3)=C(2)-K(3)\] \[C(2)-K(3)=1 \\ 2g(7,3)=6-21+C(7)-K(3) \\ 2g(7,3)=-15+C(7)-K(3) \\ 2g(7,3)-1=-15+C(7)-K(3)-C(2)+K(3) \\ 2g(7,3)=-14+C(7)-C(2) \\ g(7,3)=-7+\frac{C(7)-C(2)}{2}\]

OpenStudy (freckles):

I think I'm still making a mistake somewhere

OpenStudy (anonymous):

Oh I get it now.

OpenStudy (anonymous):

You simply use the Cauchy-Riemann equations above, and calculate the partials. There, you had \[2 = g_{y}\]

OpenStudy (anonymous):

The C-R equations are your best bet. It becomes a problem similar to finding potential functions.

OpenStudy (anonymous):

and \[3 = -g_{x}\]

OpenStudy (anonymous):

Yes, and then you find the potential function by adding the constant term.

OpenStudy (anonymous):

So your potential function g(x,y) becomes -3x + 2y + C

OpenStudy (anonymous):

But you are given an initial condition, g(2,3)=1. Plugging it back into our potential function, you will find that C = 1.

OpenStudy (anonymous):

Therefore, you get \[g(x,y) = -3x+2y+1\]

OpenStudy (anonymous):

So just plug in, g(7,3) and you will get -14.

OpenStudy (freckles):

@Loser66 check this out

OpenStudy (freckles):

I will still need to look at that theorem thing because I think I was using a little wrong. lol like I thought I was supose to write g_x=-3 implies g(x,y)=-3x+K(y) where K is a function of y and g_y=2 implies g(x,y)=2y+C(x) where C is a function of x but I kept getting ugly stuff

OpenStudy (freckles):

thanks for your help guys